# Sequential Compression Layers for Efficient Federated Learning in Foundational Models

### v1

@@ Abstract
Federated Learning (FL) has gained popularity for fine-tuning large language models (LLMs) across multiple nodes, each with its own private data. While LoRA has been widely adopted for parameter-efficient federated fine-tuning, recent theoretical and empirical studies highlight its suboptimal performance in the federated learning context. In response, we propose a novel, simple, and more effective parameter-efficient fine-tuning method that does not rely on LoRA. Our approach introduces a small multi-layer perceptron (MLP) layer between two existing MLP layers—the up_proj (the FFN projection layer following the self-attention module) and down_proj—within the feed-forward network of the transformer block. This solution addresses the bottlenecks associated with LoRA in federated fine-tuning and outperforms recent LoRA-based approaches, demonstrating superior performance for both language models and vision encoders.
@@ 1 Introduction
Large Language Models (LLMs) are central to Natural Language Processing (NLP), and fine-tuning them on task-specific datasets enhances their performance on downstream tasks. However, real-world data is often distributed and cannot be shared due to privacy concerns. Federated Learning Li et al. (2020); McMahan et al. (2017); Zhang et al. (2021) addresses this by enabling decentralized model training across institutions without exposing sensitive data. Fine-tuning foundational models like LLMs and ViTs Dosovitskiy et al. (2021) is computationally intensive, often requiring substantial resources. Parameter-efficient fine-tuning (PEFT) methods Houlsby et al. (2019) like LoRA Hu et al. (2021) help mitigate this challenge. Recently, new methods like FedFTG Mahla and Ramakrishnan (2024) have shed light on the sub-optimal nature of LoRA adapters in federated fine-tuning owing to constrained sub-spaces. Following up from this, we present a simple yet novel method of federated fine-tuning of foundational models as our main contribution of this paper. We add a small (or compressed) MLP layer between the first and second MLP layer just after the self-attention module. This new framework outperforms recent SoTA LoRA based methods like FFA-LoRA (Sun et al., 2024) and FedSA-LoRA Guo et al. (2024) on both image and text modalities.
2 Related Work
@@ 2.2 LoRA in Federated Learning
Low-Rank Adaptation (LoRA) (Hu et al., 2021) utilizes low-rank matrices to approximate gradient updates, enabling the adaptation of models without altering pre-trained weights. This method has gained popularity due to its efficiency, effectiveness, and adaptability. By significantly reducing the size of model updates, LoRA helps address communication bottlenecks in Federated Learning (FL). For instance, Yi et al. (2023) introduced FedLoRA, which integrates LoRA into FL to improve fine-tuning efficiency. Similarly, Yang et al. (2024) proposed FedDPA, a dual-adapter approach, and Qi et al. (2024) presented FDLoRA, both leveraging personalized and global LoRA modules to capture local and shared knowledge. Federated Freeze-A LoRA (FFA-LoRA) Sun et al. (2024) freezes the randomly initialized A matrices and only fine-tuning the zero-initialized B matrices, further halving communication costs and showing promising performance improvements compared to the other LoRA baselines. On the other hand, FedSA-LoRA Guo et al. (2024) which fine-tunes both matrices of low-rank adapter, but only aggregates the randomly initialized matrix A.
@@ 3 Methodology
Recent studies have provided both theoretical analysis and empirical evidence highlighting the limitations of using LoRA for federated fine-tuning of foundational models, primarily due to its constrained subspace learning Mahla and Ramakrishnan (2024). Building on these insights, we propose a novel approach to federated fine-tuning that does not rely on LoRA. Instead, our method employs direct weight averaging using FedAvg, without the need for adapters, while maintaining parameter efficiency comparable to LoRA. As illustrated in the figure, we introduce a much smaller MLP layer immediately after the MLP layer that follows the self-attention module. This results in a modification of the projection in MLP layer 2, as depicted in the figure. This additional intermediate layer is inspired by observations from Tian et al. (2024), which suggest that explicit fine-tuning of the self-attention module is unnecessary, as its information is already captured by the MLP layer directly following it. Consequently, we apply the compression layer right after the first MLP layer in the transformer block. By doing so, we manipulate the information passing through the attention module, projecting it into a smaller subspace and learning representations within this reduced dimension similar to what has been previously discussed in Mahla and Ramakrishnan (2024). During training, only the parameters of the compression layer and the new projection between compression layer and MLP layer 2 are fine-tuned.
Assuming linear activation between the projection layers, the is randomly Gaussian initialized. For any output value of the MLP-Layer 1 Projection, the following relation can be written:
@@
Here, is the gradient. Taking L2 norm on both the sides:
Theorem 1: For a convex loss , let () be the optimal LoRA parameter matrix, be the learning rate, and let the L2 norm of the gradient to be bounded (i.e. ). The excess risk () bounds for our method, involving clients and global aggregation steps having occured every local training iterations, can be expressed as follows:
Proof:
Taking L2 norm on both the sides and using the bounds of from Lemma 1:
Theorem 1 shows that the excess risk of our method has linear upper bounds, while LoRA methods such as FFA-LoRA have quadratic upper bound as shown in Mahla and Ramakrishnan (2024). In addition to this, excess risk upper bound does not depend on the number of clients .
@@ 4 Experiments
In this section, we evaluate the performance of our method across text and image modalities, focusing on its performance with LLMs and ViTs. For text, experiments involve 3 and 4 clients, while for vision, 3, 4, and 5 clients are tested. Each client holds a unique non-IID shard of a dataset.
@@ 4.1 Datasets
We conduct experiments on both text and image datasets. For text, we use the MedQuAD dataset Ben Abacha and Demner-Fushman (2019), which includes 47,457 medical question-answer pairs from 12 NIH websites, covering 39 question types. Due to MedlinePlus copyright restrictions, answers from 3 subsets were removed. We also use the Databricks Dolly-15k dataset Conover et al. (2023), which contains 15,000 high-quality human-generated prompt-response pairs for instruction tuning LLMs, with categories like brainstorming, classification, summarization, and question answering. For the vision modality, we use the Brain Tumor classification dataset Cheng (2017), comprising 3,064 T1-weighted contrast-enhanced MRI images from 233 patients, categorized into meningioma, glioma, and pituitary tumor types.
@@ 4.2 Non-IID Data Preparation
To simulate non-IID conditions, we used Dirichlet Allocation to partition each dataset into non-IID splits, following Zhang et al. (2024a); Sun et al. (2024). For text datasets (MedQuAD and Dolly-15k), we created 4 splits, and for the Brain Tumor dataset, 5 splits. The concentration parameter was set to 0.1, resulting in highly skewed distributions across clients. Splits were based on labels: question_type for MedQuAD and category for Dolly-15k. Class label distributions for MedQuAD and Dolly-15k are shown in Figure 2 and 3 respectively.
| No of clients | Method | F1 Score |
| 3 | Ours | 0.481 |
| FedSA-LoRA | 0.475 | |
| FFA-LoRA | 0.0165 | |
| 4 | Ours | 0.577 |
| FedSA-LoRA | 0.315 | |
| FFA-LoRA | 0.3175 | |
| 5 | Ours | 0.516 |
| FedSA-LoRA | 0.320 | |
| FFA-LoRA | 0.1104 |
@@
4.3 Models and Hyperparameters
For experiments on text datasets (MedQuAD and Dolly-15k), we use Gemma-2B Gemma Team et al. (2024) and Tinyllama-1.1B Zhang et al. (2024b). For our approach, the compression-layer projection weights and the new MLP-layer 2 weights were fine-tuned. In contrast, FedSA-LoRA and FFA-LoRA fine-tuned the attention modules (query, key, value) with a LoRA rank of 8 and a scaling factor of 16. For experiments on vision modality, we fine-tune SigLIP Zhai et al. (2023). For our approach, we fine-tuned the similar parameters as in language models along with the classifier layer. For FedSA-LoRA and FFA-LoRA, the attention parameters (query, key, value) were fine-tuned with a LoRA rank of 8 and a scaling factor of 32.
4.4 Experiment Results
Table 1 and 2 show results on text and image modality respectively. We train the models for 10 epochs with early stopping owing to overfitting. We report the ROUGE_L (ROUGE longest common subsequence) score and BLUE-4 (4-gram) score for text modalities while F1 score for the experiments on image modality.
As can be inferred from table 2, our method shows much better performance compared to the most recent state-of-the-art LoRA methods (FedSA-LoRA and FFA-LoRA). Similar results are observed for the image modality experiments in table 1. The notably poorer performance of LoRA methods across text and vision experiments aligns with earlier analyses, as its excess risk bounds increase with each FedAvg step (Theorem 1), hindering efficient learning of the dataset’s semantics and leading to suboptimal results.
@@ 5 Conclusion
This paper introduces a compact sequential compression layer for federated fine-tuning of foundational models, addressing limitations of LoRA-based methods like constrained subspace learning and rising excess risk. The approach projects representations into an expressive subspace, achieving superior performance with minimal parameters. Experiments on text and vision datasets demonstrate consistent improvements over state-of-the-art LoRA methods across non-IID client setups. Theoretical analysis confirms linear bounds on weight updates and excess risk, ensuring scalability and stability. This method offers an efficient, scalable solution for privacy-preserving federated learning in language and vision models, balancing computational efficiency and performance.

### v2

@@ Abstract
Federated Learning (FL) has emerged as a promising paradigm for fine-tuning large language models (LLMs) across distributed nodes, each containing private, decentralized data. While LoRA has become the de facto standard for parameter-efficient federated fine-tuning, recent theoretical and empirical analyses reveal fundamental limitations in its performance within the federated learning context. We propose a simple yet effective parameter-efficient fine-tuning method that surpasses LoRA-based approaches. Our method introduces a lightweight multi-layer perceptron (MLP) between the existing up_proj and down_proj layers within the feed-forward network of transformer blocks. Through extensive experiments, we demonstrate that our approach consistently outperforms state-of-the-art LoRA-based methods across various tasks in both language and vision modalities. While privacy considerations remain central to federated learning research, our work focuses specifically on analyzing the convergence behavior and final model performance across different federation strategies, providing empirical evidence for the better performance of our proposed architecture in distributed fine-tuning scenarios.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP), demonstrating remarkable capabilities across diverse tasks. Fine-tuning these models on domain-specific datasets has become crucial for optimizing their performance on specialized downstream tasks. However, real-world data is often distributed across institutions and cannot be centralized due to privacy concerns, regulatory requirements, and data sovereignty issues. Federated Learning Li et al. (2020); McMahan et al. (2017); Zhang et al. (2021) has emerged as a promising solution, enabling collaborative model training while keeping sensitive data localized. Fine-tuning foundation models like LLMs and Vision Transformers Dosovitskiy et al. (2021) presents significant computational challenges, requiring substantial GPU resources and memory. Parameter-efficient fine-tuning (PEFT) methods Houlsby et al. (2019), particularly Low-Rank Adaptation (LoRA) Hu et al. (2021), have gained widespread adoption by reducing the number of trainable parameters while maintaining performance. However, recent theoretical analyses and empirical studies, notably FedFTG Mahla et al. (2025), have revealed fundamental limitations in LoRA’s effectiveness within federated settings, primarily due to constrained gradient subspaces that limit model expressiveness during distributed training. Our work makes several key contributions to address these challenges:
We propose a novel parameter-efficient architecture for federated fine-tuning that introduces a compressed multi-layer perceptron (MLP) between the up-projection and down-projection layers in the transformer’s feed-forward network. This design enables more expressive parameter updates while maintaining computational efficiency.
Through extensive empirical analysis, we demonstrate that our method consistently outperforms recent approaches including FFA-LoRA (Sun et al., 2024) and FedSA-LoRA Guo et al. (2024) across both vision and language tasks. While FFA-LoRA incorporates differential privacy mechanisms, our primary focus is on evaluating their fundamental aggregation strategies and model adaptation techniques. We adopt their non-private variants (ie. we perform the experiments in the same way as was done in the experiments of the FFA-LoRA paper without considering differential privacy) as baselines to ensure a fair comparison of architectural innovations, isolating the impact of privacy mechanisms from core model performance.
We provide theoretical insights into why our approach achieves superior convergence compared to LoRA-based methods in federated settings, supported by empirical evidence across diverse tasks and data distributions.
2 Related Works
@@ 2.2 LoRA in Federated Learning
Low-Rank Adaptation (LoRA) (Hu et al., 2021) utilizes low-rank matrices to approximate gradient updates, enabling the adaptation of models without altering pre-trained weights. This method has gained popularity due to its efficiency, effectiveness, and adaptability. By significantly reducing the size of model updates, LoRA helps address communication bottlenecks in Federated Learning (FL). For instance, Yi et al. (2023) introduced FedLoRA, which integrates LoRA into FL to improve fine-tuning efficiency. Similarly, Yang et al. (2024) proposed FedDPA, a dual-adapter approach, and Qi et al. (2024) presented FDLoRA, both leveraging personalized and global LoRA modules to capture local and shared knowledge. Federated Freeze-A LoRA (FFA-LoRA) (Sun et al., 2024) freezes the randomly initialized A matrices and only fine-tuning the zero-initialized B matrices, further halving communication costs and showing promising performance improvements compared to the other LoRA baselines. FedSA-LoRA Guo et al. (2024) takes a different approach by fine-tuning both matrices but selectively aggregating only the randomly initialized matrix A. FedFMSL (Wu et al., 2024) introduces a two-stage approach combining global and local experts with a novel Sparsely Activated LoRA (SAL) algorithm. LoRA-FAIR (Bian et al., 2024) addresses two fundamental challenges in federated LoRA: server-side aggregation bias and client-side initialization drift. It introduces a correction term at the server while maintaining the original LoRA modules, demonstrating consistent performance improvements across Vision Transformer and MLP-Mixer architectures. Focusing on communication efficiency, Kuo et al. (2024) propose FLASC, which applies sparsity during communication while allowing full local LoRA fine-tuning. Their approach achieves performance comparable to dense LoRA with up to 10× reduction in communication costs, while simultaneously showing benefits for heterogeneity and privacy concerns.
@@ 3 Methodology
Recent studies have provided both theoretical analysis and empirical evidence highlighting the limitations of using LoRA for federated fine-tuning of foundational models, primarily due to its constrained subspace learning Mahla et al. (2025). Building on these insights, we propose a novel approach to federated fine-tuning that does not rely on LoRA. Instead, our method employs direct weight averaging using FedAvg, without the need for low-rank parallel adapters, while maintaining parameter efficiency comparable to LoRA. As illustrated in the figure, we introduce a much smaller MLP layer immediately after the MLP layer that follows the self-attention module. This results in a modification of the projection in MLP layer 2, as depicted in the figure. This additional intermediate layer is inspired by observations from Tian et al. (2024), which suggest that explicit fine-tuning of the self-attention module is unnecessary, as its information is already captured by the MLP layer directly following it. Consequently, we apply the compression layer right after the first MLP layer in the transformer block. By doing so, we manipulate the information passing through the attention module, projecting it into a smaller subspace and learning representations within this reduced dimension similar to what has been previously discussed in Mahla et al. (2025). During training, only the parameters of the compression layer () and the projection following it () are fine-tuned. This way, the number of parameters fine-tuned are . Clearly, having a compression layer having output dimension is analogous to having a LoRA adapter with rank but parallel to the pre-trained weight matrix showcasing the parameter efficiency that our approach brings to the table with better performance guarantees.
We randomly initialize with a unit Gaussian distribution. For any output value of the MLP-Layer 1 Projection, the following relation can be written assuming linear activation between the projection layers:
@@
Following this Lemma, we propose the following theorem.
Theorem 1:
For a convex loss , let () be the optimal parameter matrix, be the learning rate, and let the L2 norm of the gradient to be bounded (i.e. ). The excess risk () bounds for our method, involving clients and global aggregation steps having occurred every local training iterations, can be expressed as follows:
Theorem 1 demonstrates our method achieves linear excess risk bounds, independent of the number of clients , while LoRA methods have quadratic bounds that scale with Mahla et al. (2025). This advantage stems from our sequential compression design, which creates a direct path for gradient flow through a controlled subspace, preventing error accumulation during federated averaging. In contrast, LoRA’s parallel adapters can lead to divergent updates across clients, causing quadratic error accumulation.
@@ 4 Experiments
In this section, we evaluate the performance of our method across text and image modalities, focusing on its performance with LLMs and ViTs. For text, experiments involve 3 and 4 clients, while for vision, 3, 4, and 5 clients are tested. Each client holds a unique non-IID shard of a dataset created by partitioning the original dataset using Dirichlet Allocation with concentration parameter equal to 0.1.
@@ 4.1 Datasets
We conduct experiments on both text and image datasets. For text, we use the MedQuAD dataset Ben Abacha & Demner-Fushman (2019), which includes 47,457 medical question-answer pairs from 12 NIH websites, covering 39 question types. Due to MedlinePlus copyright restrictions, answers from 3 subsets were removed. We also use the Databricks Dolly-15k dataset Conover et al. (2023), which contains 15,000 high-quality human-generated prompt-response pairs for instruction tuning LLMs, with categories like brainstorming, classification, summarization, and question answering. For the vision modality, we use the Brain Tumor classification dataset Cheng (2017), comprising 3,064 T1-weighted contrast-enhanced MRI images from 233 patients, categorized into meningioma, glioma, and pituitary tumor types.
@@ 4.2 Non-IID Data Preparation
To simulate non-IID conditions, we used Dirichlet Allocation to partition each dataset into non-IID splits, following Zhang et al. (2024a); Sun et al. (2024). For text datasets (MedQuAD and Dolly-15k), we created 4 splits, and for the Brain Tumor dataset, 5 splits. The concentration parameter was set to 0.1, resulting in highly skewed distributions across clients. Splits were based on labels: question_type for MedQuAD and category for Dolly-15k. Class label distributions for MedQuAD and Dolly-15k are shown in Figure 2 and 3 respectively. A similar class label distribution for the Brain Tumor Classification dataset is shown in the appendix.
4.3 Models and Hyperparameters
For experiments on text datasets (MedQuAD and Dolly-15k), we use Gemma-2B Gemma Team et al. (2024) and Tinyllama-1.1B Zhang et al. (2024b). Following our methodology, we fine-tune two components: the compression-layer projection weights and the modified MLP-layer 2 weights. The compression layer dimension was set to 8, which maintains parameter efficiency comparable to LoRA while ensuring sufficient expressivity for learning task-specific adaptations. This choice results in trainable parameters, where is the compression dimension and are the original MLP layer dimensions. For fair comparison, we implement FedSA-LoRA and FFA-LoRA with matching parameter counts, using a LoRA rank of 8 and a scaling factor of 16 for attention module fine-tuning (query, key, value), following their original implementations. For vision experiments using SigLIP Zhai et al. (2023), we maintain architectural consistency by applying the same compression dimension () while additionally fine-tuning the classifier layer to accommodate vision-specific requirements. The baseline LoRA methods (FedSA-LoRA and FFA-LoRA) were configured with rank 8 and a scaling factor of 32, adjusted for vision tasks while maintaining comparable parameter counts to our approach.
4.4 Experiment Results
Tables 1 and 2 show results on text and image modality respectively. We train the models for 10 epochs with early stopping owing to overfitting. We report the ROUGE_L (ROUGE longest common subsequence) score and BLUE-4 (4-gram) score for text modalities while F1 score for the experiments on image modality.
@@ 4.4 Experiment Results
As demonstrated in table 1, our sequential compression layer approach significantly outperforms recent state-of-the-art LoRA methods (FedSA-LoRA and FFA-LoRA) across text-based tasks. Similar performance gains are observed in vision modality experiments (table 2). These results empirically validate our theoretical analysis - while LoRA-based methods suffer from quadratic excess risk bounds that scale with both the number of clients and aggregation steps, our method’s linear bounds () enable more stable learning. The strategic placement of our compression layer between MLP layers, rather than parallel adapters as in LoRA, allows for more expressive parameter updates within a controlled subspace, leading to better capture of task-specific semantics across both modalities while maintaining comparable parameter efficiency.
| No of clients | Method | F1 Score |
| 3 | Ours | 0.481 |
| FedSA-LoRA | 0.475 | |
| FFA-LoRA | 0.0165 | |
| 4 | Ours | 0.577 |
| FedSA-LoRA | 0.315 | |
| FFA-LoRA | 0.3175 | |
| 5 | Ours | 0.516 |
| FedSA-LoRA | 0.320 | |
| FFA-LoRA | 0.1104 |
@@ 5 Conclusion
This paper presents a novel approach to federated fine-tuning of foundational models through strategically placed sequential compression layers, offering both theoretical and practical advantages over existing LoRA-based methods. Our key innovation lies in projecting representations into a carefully designed compressed subspace immediately after the self-attention module, enabling more expressive parameter updates while maintaining efficiency. Unlike LoRA-based approaches that suffer from constrained subspace learning and quadratic excess risk bounds, our method achieves linear bounds on both weight updates and excess risk, independent of the number of clients. This theoretical advantage translates to practical benefits, as demonstrated through extensive experiments across both language (Gemma-2B, TinyLlama-1.1B) and vision (SigLIP) models under challenging non-IID data distributions. The empirical results consistently show superior performance over state-of-the-art methods like FFA-LoRA and FedSA-LoRA while maintaining comparable parameter efficiency. Our approach’s ability to achieve better results makes it particularly suitable for real-world federated learning deployments where resources are constrained and data privacy is crucial. Furthermore, the method’s simplicity and theoretical guarantees provide a solid foundation for future research in efficient, privacy-preserving distributed learning of foundation models.
Appendix A Appendix
A.1 Proofs
where is the learning rate and B is a constant.
Here, is the gradient. Taking L2 norm on both the sides:
Theorem 1: For a convex loss , let () be the optimal parameter matrix, be the learning rate, and let the L2 norm of the gradient to be bounded (i.e. ). The excess risk () bounds for our method, involving clients and global aggregation steps having occurred every local training iterations, can be expressed as follows:
Proof:
Taking L2 norm on both the sides and using the bounds of from Lemma 1:
A.2 Dataset Shard Class Distribution
Below figure shows the class label distributions for the Brain Tumor Classification Dataset. The dataset is sharded using Dirichlet Allocation with a concentration parameter .